OpenStudy (anonymous):

point c is the midpoint of line segment AB. Point A has coordinates (4,8) and point C has coordinates (-8,12). what are the coordinates of point B?

OpenStudy (anonymous):

If B has coordinates (x, y), then \[ \frac {4+x}{2} = -8, \qquad \frac {8+y}{2} = 12. \]Can you finish now?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

To solve the first equation: \[ \frac {4+x}{2} = -8 \implies 4 + x = -16 \implies x = -20. \]Can you do the second one now?